OpenStudy (anonymous):

I have a few questions: They are equations with variables on each side" 1) 2h-8=h+17 2) 7a-3=3-2a 3) -6y-3=3-6y 4) 2(3u+7= -4(3-2u) P.S ] please tell me how to get the answer. Thank you very much beforehand.

OpenStudy (shamil98):

2h - 8 = h + 17 all you have to do is get h on one side and everything else on the other side.. 2h - 8 = h + 17 - h - h isolate h h - 8 = 17 and solve for h from there. + 8 +8 h = 25

OpenStudy (anonymous):

thank you! If you don't have a problem could help me with the others

OpenStudy (shamil98):

I've just explained how to do it, try doing the next one on your own.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

but if I need help then I'm going to come back so don't stay waiting I'm going to try and do.

OpenStudy (shamil98):

okay.

OpenStudy (anonymous):

okay so here are my answers 2. 1.5 3. 0.5 4. -1.8